1. A company will have to spend $300000 on new plant in two years from now.

Currently investment rates are at a nominal 10%. (a) What single sum should now be invested, if compounding is six-monthly?


(b) What is the APR?



  1. Find the effective rate (APR) for a nominal rate at


(a) 7% p.a. compounded annually.


(b) 8% p.a. compounded semi-annually.


(c) 10% p.a. compounded quarterly.


(d) 9% p.a. compounded monthly.


1
Expert's answer
2021-11-19T01:26:25-0500

1a.

Present Value = $300,000

Time Period = 2 Years

Interest Rate = 10% 


"Future \\space Value = Present\\space \u2009Value \u00d7(1+Interest\\space Rate)^{Time\\space Period}\\\\ = \\$300,000 \u00d7(1+\\frac{0.1}{2})^{2\u00d72}\\\\ = \\$300,000 \u00d7(1+0.05)^{4}\\\\ = \\$300,000 \u00d7(1.05)^4\\\\ = \\$300,000 \u00d71.21550625 \\\\ = \\$364,651.875 \\space or\\space \\$364,651.88"


b.

"Annual \space Percentage\space Rate\\ =\frac{2×Time Period×Interest rate}{Time Period +1}\\ =\frac{2×2×10\%}{2 +1}\\ =\frac{0.4}{3}\\ =0.1333333333 \space or \space 13.33\%"

1.

Answer A.

Under annual compounding, the number of compounding periods will be equal to 1.

Effective Annual Rate = Nominal Rate = 7%

Answer B.

Nominal Rate = 8%

Compounded Semi Annually

Number of compounding periods = 2

Effective Interest Rate =(1+(Nominal Rate"\div" Number of compounding periods))

Number of compounding periods − 1

"= (1+(\frac{8\%}{2}))^2 − 1\\
=1.0816− 1\\
 = 8.16\%"

Answer C.

Nominal Rate = 10%

Compounded Quarterly

Number of compounding periods = 4

Effective Interest Rate =(1+(Nominal Rate"\div" Number of compounding periods))

Number of compounding periods − 1

"= (1+(\frac{10\%}{4}))^4 − 1\\
=1.1038− 1\\
 = 10.38\%"

Answer D.

Nominal Rate = 9%

Compounded monthly

Number of compounding periods = 12

Effective Interest Rate =(1+(Nominal Rate"\div" Number of compounding periods))

Number of compounding periods − 1

"= (1+(\frac{9\%}{12}))^{12} − 1\\
=1.0938− 1\\
 = 9.38\%"